# Monte Carlo Physics-informed neural networks for multiscale heat conduction via phonon Boltzmann transport equation

### v2

@@ Abstract
The phonon Boltzmann transport equation (BTE) is widely used for the description of multiscale heat conduction (from $nm$ to $\mu m$ or $mm$) in solid materials. Developing numerical approaches to solve this equation is challenging since it is a 7-dimensional integral-differential equation. In this work, we propose the Monte Carlo physics-informed neural networks (MC-PINNs), which do not suffer from the “curse of dimensionality”, to solve the phonon Boltzmann transport equation to model the multiscale heat conduction in solid materials. In the MC-PINNs, we utilize a deep neural network to approximate the solution to the BTE, and encode the BTE as well as the corresponding boundary/initial conditions using the automatic differentiation. In addition, we propose a novel two-step sampling approach to address the issues of inefficiency and the inaccuracy issues in the widely used sampling methods in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial space (Step I), and then draw another number of points randomly in the solid angular space (Step II). The training points are constructed given the data drawn from the above two steps using the tensor product. The two-step sampling strategy enables the MC-PINNs (1) to model the heat conduction from ballistic to diffusive regimes, and (2) is memory more efficient compared to the conventional numerical solvers or existing PINNs for BTE. A series of numerical examples including a quasi-one-dimensional (quasi-1D) steady/unsteady heat conduction in a film, and the heat conduction in a quasi-two- and three-dimensional square domain, are conducted to justify the effectiveness of the MC-PINN for heat conduction spanning diffusive and ballistic regimes. Finally, we perform a comparison on the computational time and the memory usage of the MC-PINN and one of the state-of-art numerical methods to demonstrate the potential of the MC-PINN for large scale problems in real-world applications.
@@ 1 Introduction
Microdevices, e.g., microprocessors and microelectronic circuits, often generate heat with great density during operation. Efficient thermal management is crucial to the reliability and the optimal performance of these devices. As pointed out in [1], the heat conduction in microdevices like chips is inherently a multiscale problem spanning nanoscale (1-100 nm) to microscale (0.1-100 $\mu$m) or macroscopic (0.1-100 mm). The classical Fourier’s law is a widely used empirical macroscopic model for heat conduction in solids. However, it is not capable of describing the heat conduction in the non-diffusive or ballistic regime [2, 3, 4]. Further, the phonon Boltzmann transport equation (BTE) is generally employed for describing the heat conduction in solid materials spanning the ballistic and diffusive regimes [5, 1, 6, 7].
To advance the understanding of multiscale heat conduction in solid materials, lots of efforts have been put into developing numerical methods for solving the phonon BTE. As mentioned, developing effective numerical methods for solving the phonon BTE efficiently is challenging since it is a 7-dimensional integral differential equation in real-world applications. The conventional numerical methods can be broadly categorized into two main approaches, i.e., the stochastic methods, e.g., direct simulation Monte Carlo (DSMC) [8, 9, 10, 11], etc., and deterministic approaches, e.g., discrete velocity/ordinate method [12, 13, 14], synthetic iterative scheme [15, 16, 17], discrete unified gas kinetic scheme (DUGKS) [18, 19], and so on. The former uses discrete meshes in the temporal-spatial space and is efficient for steady problems in the ballistic regime. However, it converges slowly for problems in the (near) diffusive regime as well as non-stationary problems. The latter employs the discrete meshes in the temporal-spatial-angular space and is able to achieve good accuracy for steady/unsteady problems spanning a wide range of regimes. In general, the number of discrete points in the solid angular space required to obtain accurate results for the heat conduction problem in the diffusive regime is much smaller than in the ballistic regime. The requirement of large numbers of discrete points in the solid angular space for accurately capturing the non-Fourier’s behavior in the ballistic regime leads to intensive computational cost and prohibitive memory usage for large-scale thermal engineering problems in practice.
@@ 2.1 Phonon Boltzmann transport equation
The phonon Boltzmann transport equation (BTE) is a fundamental tool for understanding heat conduction in solid materials at different scales. Various models of the phonon BTE have been developed to address specific physical scenarios [9]. For instance, the full/linearized phonon BTE considers the detailed phonon dispersion and full scattering kernel, enabling the study of complex interactions among different phonon modes [30]. For scenarios that demand frequency-specific considerations, the frequency-resolved BTE under relaxation time approximation is often employed, in which the relaxation rate on the diagonal of phonon scattering matrix is accounted [10, 11, 16, 17]. A widely used model among these approaches is the gray model, where nonlinear phonon dispersion and polarization are not accounted for and frequency-independent assumption is used [19, 15, 17]. In this study, we focus on the phonon gray model, which is apt for probing multiscale heat conduction in three-dimensional materials subject to minor temperature variations, i.e., $\Delta T\ll T_{ref}$. Based on a certain number of reference variables, e.g., reference temperature $T_{ref}$, specific heat $C_{ref}$, reference length $L_{ref}$, group velocity $v_{g}$ and time $t_{ref}=L_{ref}/v_{g}$, then the dimensionless phonon BTE in the gray model can be obtained as
@@
We would like to discuss that the multiscale heat conduction is generally divided into different regimes based on the Knudsen number, seen in Figure 1. When $\rm{Kn}<0.01$, the characteristic length of the system is much larger than the phonon mean free path and phonon transport is in the diffusive regime. The resulting temperature field is linear, consistent with the predictions of Fourier’s law of heat conduction. As $\rm{Kn}$ increases, Fourier’s law begins to fail. When $\rm{Kn}>10$, heat transfer is dominated by ballistic effects.
@@ 2.2.1 Physics-informed neural networks
As for the selection of the training points for evaluating the loss function in PINNs (Eq. (7)), two different ways are often used, i.e., (1) the number ($N/N_{BC}/N_{IC}$) as well as the locations of the training points are prescribed at each training step, and (2) the number of points ($N/N_{BC}/N_{IC}$) is prescribed, but the corresponding locations are randomly generated at each training step. The former is similar to the meshes in the conventional numerical methods. As mentioned, lots of discrete points in the solid angular space are required to accurately capture the non-Fourier behavior at the ballistic regime, which results in expensive computational and memory cost for multiscale heat conduction problems in real-world applications. For the latter approach, if we randomly generate the training points in the temporal-spatial-angular domain, denoted as $(t,\bm{x},\bm{s})_{i},i=1,...,M$, it is probable that we do not have two points that are exactly the same, i.e., $(t,\bm{x},\bm{s})_{i}$ are different for different $i$. In other words, we only update the energy density function at $(t,\bm{x})_{i}$ corresponding to one angular point, i.e., $\bm{s}_{i}$. It leads to slow convergence or even degenerated results since we need to compute the integral in the entire angular domain, as we will show in Sec. 3.1.
@@ 2.2.2 Two-step sampling approach
Here we further discuss the superiority of the two-step sampling approach over the the first sampling method in Sec. 2.2.1. As aforementioned, a large number of points in the solid angular domain has to be employed in the ballistic regime for the first approach of Sec. 2.2.1. In addition, how to determine the optimal number of discrete points at different regimes remains an open question. In the two-step sampling method, we randomly generate a certain number of the points in the solid angular space at each training step. We can then traverse the entire angular domain even we employ a small number of points in each step as long as we have enough training steps, which is similar to the statistic or particle methods for solving the BTE. Hence, compared to the discrete ordinate/velocity methods, the present solver is free of ray effects in the ballistic regime. Therefore, less computational and memory cost for heat conduction at the ballistic regimes is then expected at each training step if the two-step sampling strategy is adopted. Also, as we will show in Sec. 3.1, the two-step sampling method is more flexible than the first sampling approach in Sec. 2.2.1 for multiscale heat conduction problems. For the implementations of boundary/initial conditions, we can randomly generate a certain number of training points in the temporal-spatial-angular domain, i.e., $B_{t,\bm{x},\bm{s}}$, since we do not need to compute the integral in boundary/initial conditions, for computational efficiency.
@@ 3 Results and Discussion
In this section, we test the proposed method for simulating the heat conduction in solids ranging from diffusive to ballistic regimes, using four demonstration examples: (1) steady and unsteady quasi-one-dimensional (quasi-1D) film heat conduction, (2) quasi-two-dimensional (quasi-2D) heat conduction in a square domain, and (3) three-dimensional (3D) cuboid phonon transport. Further, the Adam optimizer, which is a variant of the stochastic gradient descent method, is utilized to minimize the loss functions in MC-PINNs. The details for computations, e.g., the architectures of NNs, activation functions, and the numbers of training steps, etc., as well as the details for reference solutions are present in A.
@@ 3.1 Quasi-1D film heat conduction
Figure 3: Schematic of the quasi-one-dimensional film heat conduction problem. $T_{h}$ and $T_{c}$ are the temperatures at the left and right boundaries, respectively; and $L$ is the thickness of the film.
@@
Figure 5: Quasi-1D film heat conduction: Predicted (a) temperature, and (b) heat flux, from MC-PINNs at different Knudsen numbers. Black solid line: MC-PINNs; Symbols: Reference solution [16]. Gray dashed line: the solution of Fourier’s Law.
@@
Figure 6: Quasi-1D film heat conduction: Effect of different sampling approaches in solid angular space on the predicted accuracy at (a) $\rm{Kn}=0.01$, (b) $\rm{Kn}=1$, and (c) $\rm{Kn}=10$. Symbols: Reference solution; Line 1: Random sampling in both physical and solid angular space; Line 2: uniform discrete points in both physical and solid angular space; Line 3: uniform discrete points in physical space and GL quadrature points; Line 4: uniform discrete points in physical space and random sampling in solid angular space; Line 5: random sampling in physical-angular space.
We utilize these five sampling methods in PINNs to model the heat conduction in the 1D film at $\rm{Kn}=0.01$, 1, and 10. The training step is set as 50,000 at $\rm{Kn}=1,10$ and 10,$0$ at $\rm{Kn}=0.01$ in all test cases. The results and the computational errors are present in Fig. 6 and Table 2, respectively. As we can see, (1) PINNs with the two-step sampling approach, i.e., PINNs-1 and PINN-4, is able to provide the most accurate results for the heat conduction from diffusive to ballistic regimes; (2) the method in [27], i.e., PINNs-3, is able to achieve good accuracy for Kn $=0.01$ and 1, but failed to capture the non-Fourier’s behavior accurately at Kn $=10$ (e.g., the temperature jumps at the boundaries), as shown in Fig. 6; and (3) the computational errors for PINN-5 increases with the decreasing Kn. As discussed in Sec. 2.2.2, we only solve the BTE at one point in the solid angular space at each spatial point, leading to the poor accuracy for the computation of $e^{eq}$. At the ballistic scale, i.e., cases with large Kn, we have very few collision for the phonon, and the inaccurate estimation of $e^{eq}$ has little effect on the predicted accuracy. However, for cases with small Kn, we have frequent collision of the phonon and the inaccurate prediction of $e^{eq}$ results in poor accuracy for solving the BTE.
@@ 3.2 Unsteady quasi-1D film heat conduction with varying Knudsen number
where $\rm{Kn}_{\max}$ and $\rm{Kn}_{\min}$ represent the maximum and minimum Knudsen number, respectively; $x_{c}=L/2$ is the center of the film with $L$ denoting the thickness of the film, and $d=0.01$ characterizes the thickness of the transition layer. In the present case, we set $\rm{Kn}_{\max}=10$ and $\rm{Kn}_{\min}=0.01$. The distribution of the Knudsen number within the film is shown in Fig. 7. It can be seen that we have a significant change in $\rm{Kn}$ near the center of the film. The heat conduction at the left half is at the ballistic regime, which it is at the diffusive regime on the right half of the film.
@@
Figure 8: Unsteady quasi-1D film heat conduction: The predicted temperature at different times with space varying Knudsen number. (a) $t=0.2$, (b) $t=0.5$, (c) $t=1.0$, (d) $t=10.0$.
@@ 3.3 Quasi-2D square heat conduction
The results of MC-PINNs are shown in Figs. 10 and 11. We can see that the predicted temperatures from the MC-PINNs are in good agreement with the reference solutions in all test cases. We also observe in Figs. 10 and 11 there are significant temperature jumps at the left/right boundaries. In addition, the temperature jump becomes more pronounced with the increase of $\rm{Kn}$. This phenomenon is reasonable since the non-equilibrium effects that become more pronounced in the transition and ballistic regimes, leading to distinct temperature discontinuities at the walls.
Figure 10: Quasi-2D heat conduction: Distribution of the temperature at (a) $\rm{Kn}=0.1$, (b) $\rm{Kn}=1.0$ and (c) $\rm{Kn}=10.0$. Colored background with black solid line: reference solution [17]; red dash line: the MC-PINNs.
@@
The predicted temperature from MC-PINNs in each case at $x=0.5$ are present in Figs. 13 and 14. Again, good agreements are observed for all test cases between the MC-PINNs and the reference solutions. Similar as in the 2D cases, there are significant temperature jumps at the boundaries when $\rm{Kn}\geq 1$, which have been discussed in Sec. 3.3 and will not present here for simplicity. We then illustrate the relative $L_{2}$ errors between the temperature from the MC-PINNs and the reference solutions at different $\rm{Kn}$ in Table 4. The results further confirm the capability of the MC-PINNs for 3D multiscale heat conduction problems.
Figure 13: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5$ for different $\rm{Kn}$. Colored background with black solid line: reference solution; Red dash line: the MC-PINNs.
Figure 14: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5,y=0.5$. Red solid line: MC-PINNs; Symbbols: Reference solution.
@@
As mentioned, the method in [17] is one of the most computationally and memory efficient approaches for solving stationary BTE. We now perform a comparison on the computational and the memory cost between the MC-PINNs and the numerical method [17] that is used to generate the reference solution here. Specifically, we compare the computational time and the memory usage using two representative cases, i.e., Kn $=0.1$ and 10.
We first discuss the computational time for MC-PINNs. As illustrated in Table 5, the computational time of MC-PINNs (second column in Table 5) is about 7 and 1.3 times of that in the numerical method [17] for Kn $=0.1$ and 10, respectively. In the deep learning community, the transfer learning technique is a widely used approach to speed up the training of DNNs. Here we can also use it to speedup the training of MC-PINNs. Specifically, we first train the MC-PINNs from scratch for Kn $=1$, and then save the hyperparamters, i.e., weights and biases, in order to reuse them in other cases. In particular, we initialize the MC-PINNs for Kn $=0.1$ and 10 using the hyperparameters from the pretrained MC-PINNs for Kn $=1$, and further train them for an additional number of steps until the convergence is reached. The relative $L_{2}$ errors for these two cases with transfer learning are $2.40\%$ and $2.6\%$, which are similar as the cases without transfer learning. Also, as we can see Table 5, we can achieve about 3-6 times speedup with transfer learning, and the computational time for these two cases are now comparable to the numerical method in [17].
Note that we run the numerical method and MC-PINNs on CPU and GPU, respectively. In particular, all the CPU computations are conducted on a single Intel Xeon Platinum 9242 processor, and the GPU computations are performed using an NVIDIA GeForce RTX 3090. We remark that the comparison present in Table 5 is not fair since we can also accelerate the numerical method by developing GPU codes and run the tests on GPUs. However, developing GPU codes for the numerical method is more challenging than MC-PINNs and is beyond the scope of the current study. We show the results here to provide a preliminary reference on the computational cost of MC-PINNs and one of the state-of-art numerical methods in this field. Therefore, we only present the computational time for the numerical method on CPUs.
@@
We now discuss the memory usage for MC-PINNs and the numerical method, which is present in Table 6. Note that the MC-PINNs with and without transfer learning has the same memory cost. We therefore only illustrate the memory usage in the MC-PINNs without transfer learning in Table 6. In the MC-PINNs, we employ the same number of training points for Kn $=0.1$ and 10, and thus the memory usage is almost the same for these two cases. In the numerical method, the memory cost is more expensive for Kn $=10$ than for Kn $=0.1$ since we have more discrete points in the angular domain for the former case. Also, the MC-PINN requires about $66\%$ and $16\%$ of the memory cost in the numerical method [17] for Kn $=0.1$ and 10, respectively. It is worth mentioning that the problem we consider here is relatively small. For real-world applications, we may have more than 40 meshes in each direction of the spatial domain in numerical methods. For example, if we increase the number of meshes in each direction from 40 to 400 and keep the same discretization in the solid angular space for Kn $=10$, the memory usage will increase to about 1,500GB, which is quite challenging to handle. However, in the MC-PINNs, we can use the same setup as used in the present case, suggesting that the memory usage can be the same, i.e., which is about 6,000 times less than in the numerical method [17].
@@
We would like to briefly summarize our main findings here: (1) the two-step sampling approach enables the MC-PINNs to use much less memory compared to the numerical method [17], especially for large scale heat conduction problems spanning the diffusive and ballistic regimes; and (2) the computational time of MC-PINNs (GPU time) with transfer learning is comparable to the numerical method [17] (CPU time) for the specific problem considered here.

### v3

@@ Abstract
The phonon Boltzmann transport equation (BTE) is widely used for the description of multiscale heat conduction (from $nm$ to $\mu m$ or $mm$) in solid materials. Developing numerical approaches to solve this equation is challenging since it is a 7-dimensional integral-differential equation. In this work, we propose the Monte Carlo physics-informed neural networks (MC-PINNs), which do not suffer from the “curse of dimensionality”, to solve the phonon Boltzmann transport equation for modeling the multiscale heat conduction in solid materials. In MC-PINNs, we utilize a deep neural network to approximate the solution to the BTE and encode the BTE as well as the corresponding boundary/initial conditions using the automatic differentiation. In addition, we propose a novel two-step sampling approach to address the issues of inefficiency and inaccuracy issues in the widely used sampling methods in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial space (Step I) and then draw another number of points randomly in the solid angular space (Step II). The training points are constructed based on the data drawn from the above two steps using the tensor product. The two-step sampling strategy enables the MC-PINNs (1) to model the heat conduction from ballistic to diffusive regimes, and (2) is memory more efficient compared to the conventional numerical solvers or existing PINNs for BTE. A series of numerical examples including a quasi-one-dimensional (quasi-1D) steady/unsteady heat conduction in a film, and the heat conduction in a quasi-two- and three-dimensional square domain, are conducted to justify the effectiveness of the MC-PINN for heat conduction spanning diffusive and ballistic regimes. Finally, we perform a comparison of the computational time and the memory usage of the MC-PINN with one of the state-of-art numerical methods to demonstrate the potential of the MC-PINN for large-scale problems in real-world applications.
@@ 1 Introduction
Microdevices, e.g., microprocessors and microelectronic circuits, often generate heat with great density during operation. Efficient thermal management is crucial to the reliability and optimal performance of these devices. As pointed out in [1], the heat conduction in microdevices like chips is inherently a multiscale problem spanning nanoscale (1-100 nm) to microscale (0.1-100 $\mu$m) or macroscopic (0.1-100 mm). The classical Fourier’s law is a widely used empirical macroscopic model for heat conduction in solids. However, it is not capable of describing the heat conduction in the non-diffusive or ballistic regime [2, 3, 4]. Further, the phonon Boltzmann transport equation (BTE) is generally employed for describing the heat conduction in solid materials spanning the ballistic and diffusive regimes [5, 1, 6, 7].
To advance the understanding of multiscale heat conduction in solid materials, lots of efforts have been put into developing numerical methods for solving the phonon BTE. As mentioned, developing effective numerical methods for efficiently solving the phonon BTE is challenging since it is a 7-dimensional integral differential equation in real-world applications. The conventional numerical methods can be broadly categorized into two main approaches, i.e., the stochastic methods, e.g., direct simulation Monte Carlo (DSMC) [8, 9, 10, 11], etc., and deterministic approaches, e.g., discrete velocity/ordinate method [12, 13, 14], synthetic iterative scheme [15, 16, 17], discrete unified gas kinetic scheme (DUGKS) [18, 19], and so on. The former uses discrete meshes in the temporal-spatial space and is efficient for steady problems in the ballistic regime. However, it converges slowly for problems in the (near) diffusive regime as well as non-stationary problems. The latter employs the discrete meshes in the temporal-spatial-angular space and is able to achieve good accuracy for steady/unsteady problems spanning a wide range of regimes. In general, the number of discrete points in the solid angular space required to obtain accurate results for the heat conduction problem in the diffusive regime is much smaller than in the ballistic regime. The requirement of large numbers of discrete points in the solid angular space for accurately capturing the non-Fourier’s behavior in the ballistic regime leads to intensive computational cost and prohibitive memory usage for large-scale thermal engineering problems in practice.
@@ 2.1 Phonon Boltzmann transport equation
The phonon Boltzmann transport equation (BTE) is a fundamental tool for understanding heat conduction in solid materials at different scales. Various models of the phonon BTE have been developed to address specific physical scenarios [9]. For instance, the full/linearized phonon BTE considers the detailed phonon dispersion and full scattering kernel, enabling the study of complex interactions among different phonon modes [30]. For scenarios that demand frequency-specific considerations, the frequency-resolved BTE under relaxation time approximation is often employed, in which the relaxation rate on the diagonal of phonon scattering matrix is accounted [10, 11, 16, 17]. A widely used model among these approaches is the gray model, where nonlinear phonon dispersion and polarization are not accounted for and the frequency-independent assumption is used [19, 15, 17]. In this study, we focus on the phonon gray model, which is apt for probing multiscale heat conduction in three-dimensional materials subject to minor temperature variations, i.e., $\Delta T\ll T_{ref}$. Based on a certain number of reference variables, e.g., reference temperature $T_{ref}$, specific heat $C_{ref}$, reference length $L_{ref}$, group velocity $v_{g}$ and time $t_{ref}=L_{ref}/v_{g}$, then the dimensionless phonon BTE in the gray model can be obtained as
@@
We would like to discuss that the multiscale heat conduction is generally divided into different regimes based on the Knudsen number, seen in Figure 1. When $\rm{Kn}<0.01$, the characteristic length of the system is much larger than the phonon mean free path, and phonon transport is in the diffusive regime. The resulting temperature field is linear, consistent with the predictions of Fourier’s law of heat conduction. As $\rm{Kn}$ increases, Fourier’s law begins to fail. When $\rm{Kn}>10$, heat transfer is dominated by ballistic effects.
@@ 2.2.1 Physics-informed neural networks
As for the selection of the training points for evaluating the loss function in PINNs (Eq. (7)), two different ways are often used, i.e., (1) the number ($N/N_{BC}/N_{IC}$) as well as the locations of the training points are prescribed at each training step, and (2) the number of points ($N/N_{BC}/N_{IC}$) is prescribed, but the corresponding locations are randomly generated at each training step. The former is similar to the meshes in the conventional numerical methods. As mentioned, lots of discrete points in the solid angular space are required to accurately capture the non-Fourier behavior at the ballistic regime, which results in expensive computational and memory costs for multiscale heat conduction problems in real-world applications. For the latter approach, if we randomly generate training points in the temporal-spatial-angular domain, denoted as $(t,\bm{x},\bm{s})_{i},i=1,...,M$, it is likely that each point $(t,\bm{x},\bm{s})_{i}$ will be unique, meaning that for different $i$, the points $(t,\bm{x})_{i}$ correspond to different angular direction $\bm{s})_{i}$. As a result, during the training process, we only update the energy density function at each $(t,\bm{x})_{i}$ based on one single angular direction $\bm{s}_{i}$. It can lead to slow convergence or even degenerated results since we need to compute the integral in the entire angular domain, as we will show in Sec. 3.1.
@@ 2.2.2 Two-step sampling approach
Here we further discuss the superiority of the two-step sampling approach over the the first sampling method in Sec. 2.2.1. As aforementioned, a large number of points in the solid angular domain has to be employed in the ballistic regime for the first approach of Sec. 2.2.1. In addition, how to determine the optimal number of discrete points at different regimes remains an open question. In the two-step sampling method, we randomly generate a certain number of points in the solid angular space at each training step. We can then traverse the entire angular domain even if we employ a small number of points in each step as long as we have enough training steps, which is similar to the statistic or particle methods for solving the BTE. Hence, compared to the discrete ordinate/velocity methods, the present solver is free of ray effects in the ballistic regime. Therefore, less computational and memory cost for heat conduction at the ballistic regimes are then expected at each training step if the two-step sampling strategy is adopted. Also, as we will show in Sec. 3.1, the two-step sampling method is more flexible than the first sampling approach in Sec. 2.2.1 for multiscale heat conduction problems. For the implementations of boundary/initial conditions, we can randomly generate a certain number of training points in the temporal-spatial-angular domain, i.e., $B_{t,\bm{x},\bm{s}}$, since we do not need to compute the integral in boundary/initial conditions, for computational efficiency.
@@ 3 Results and Discussion
In this section, we test the proposed method for simulating the heat conduction in solids ranging from diffusive to ballistic regimes, using four demonstration examples: (1) steady and unsteady quasi-one-dimensional (quasi-1D) film heat conduction, (2) quasi-two-dimensional (quasi-2D) heat conduction in a square domain, and (3) three-dimensional (3D) cuboid phonon transport. Further, the Adam optimizer, which is a variant of the stochastic gradient descent method, is utilized to minimize the loss functions in MC-PINNs. The details for computations, e.g., the architectures of NNs, activation functions, and the numbers of training steps, etc., as well as the details for reference solutions, are present in A.
@@ 3.1 Quasi-1D film heat conduction
Figure 3: Schematic of the quasi-1D film heat conduction problem. $T_{h}$ and $T_{c}$ are the temperatures at the left and right boundaries, respectively; and $L$ is the thickness of the film.
@@
Figure 5: Quasi-1D film heat conduction: Predicted (a) temperature, and (b) heat flux, from MC-PINNs at different Knudsen numbers. Solid line: MC-PINNs; Symbols: Reference solution [16]. Gray dashed line: the solution of Fourier’s Law.
@@
Figure 6: Quasi-1D film heat conduction: Effect of different sampling approaches in solid angular space on the predicted accuracy at (a) $\rm{Kn}=0.01$, (b) $\rm{Kn}=1$, and (c) $\rm{Kn}=10$. Symbols: Reference solution; PINNs-1: Two-step sampling method: Random sampling in both spatial and solid angular domain; PINNs-2: uniform discrete points in both spatial and solid angular domain; PINNs-3: uniform discrete points in spatial domain and GL quadrature points in angular domain; PINNs-4: Two-step sampling method: uniform discrete points in spatial domain and random sampling in solid angular space; PINNs-5: random sampling in spatial-angular domain.
We utilize these five sampling methods in PINNs to model the heat conduction in the 1D film at $\rm{Kn}=0.01$, 1, and 10. The training step is set as 50,000 at $\rm{Kn}=1,10$ and 10,$0$ at $\rm{Kn}=0.01$ in all test cases. The results and the computational errors are present in Fig. 6 and Table 2, respectively. As we can see, (1) PINNs with the two-step sampling approach, i.e., PINNs-1 and PINN-4, are able to provide the most accurate results for the heat conduction from diffusive to ballistic regimes; (2) the method in [27], i.e., PINNs-3, is able to achieve good accuracy for Kn $=0.01$ and 1, but failed to capture the non-Fourier’s behavior accurately at Kn $=10$ (e.g., the temperature jumps at the boundaries), as shown in Fig. 6; and (3) the computational errors for PINN-5 increases with the decreasing Kn. As discussed in Sec. 2.2.2, we only solve the BTE at one point in the solid angular space at each spatial point, leading to poor accuracy for the computation of $e^{eq}$. At the ballistic scale, i.e., cases with large Kn, we have very few collisions for the phonon, and the inaccurate estimation of $e^{eq}$ has little effect on the predicted accuracy. However, for cases with small Kn, we have frequent collisions of the phonon, and the inaccurate prediction of $e^{eq}$ results in poor accuracy for solving the BTE.
@@ 3.2 Unsteady quasi-1D film heat conduction with varying Knudsen number
where $\rm{Kn}_{\max}$ and $\rm{Kn}_{\min}$ represent the maximum and minimum Knudsen number, respectively; $x_{c}=L/2$ is the center of the film with $L$ denoting the thickness of the film, and $d=0.01$ characterizes the thickness of the transition layer. In the present case, we set $\rm{Kn}_{\max}=10$ and $\rm{Kn}_{\min}=0.01$. The distribution of the Knudsen number within the film is shown in Fig. 7. It can be seen that we have a significant change in $\rm{Kn}$ near the center of the film. The heat conduction at the left half is at the ballistic regime, which is at the diffusive regime on the right half of the film.
@@
Figure 8: Unsteady quasi-1D film heat conduction: The predicted temperature at different times with space varying Knudsen number. (a) $t=0.2$, (b) $t=0.5$, (c) $t=1.0$, (d) $t=10.0$. Symbols: Reference solution [19]; Black solid line: MC-PINNs.
@@ 3.3 Quasi-2D square heat conduction
The results of MC-PINNs are shown in Figs. 10 and 11. We can see that the predicted temperatures from the MC-PINNs are in good agreement with the reference solutions in all test cases. We also observe in Figs. 10 and 11 there are significant temperature jumps at the left/right boundaries. In addition, the temperature jump becomes more pronounced with the increase of $\rm{Kn}$. This phenomenon is reasonable since the non-equilibrium effects become more pronounced in the transition and ballistic regimes, leading to distinct temperature discontinuities at the walls.
Figure 10: Quasi-2D heat conduction: Distribution of the temperature at (a) $\rm{Kn}=0.1$, (b) $\rm{Kn}=1.0$ and (c) $\rm{Kn}=10.0$. Colored background with black solid line: reference solution [17]; Red dashed line: the MC-PINNs.
@@
The predicted temperature from MC-PINNs in each case at $x=0.5$ are present in Figs. 13 and 14. Again, good agreements are observed for all test cases between the MC-PINNs and the reference solutions. Similar to in the 2D cases, there are significant temperature jumps at the boundaries when $\rm{Kn}\geq 1$, which have been discussed in Sec. 3.3 and will not present here for simplicity. We then illustrate the relative $L_{2}$ errors between the temperature from the MC-PINNs and the reference solutions at different $\rm{Kn}$ in Table 4. The results further confirm the capability of the MC-PINNs for 3D multiscale heat conduction problems.
Figure 13: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5$ for different $\rm{Kn}$. Colored background with black solid line: reference solution; Red dashed line: the MC-PINNs.
Figure 14: 3D heat conduction problem: Predicted temperature from MC-PINNs at $x=0.5,y=0.5$. Solid line: MC-PINNs; Symbols: Reference solution [17].
@@
As mentioned, the method in [17] is one of the most computationally and memory-efficient approaches for solving stationary BTE. We now perform a comparison of the computational and the memory cost between the MC-PINNs and the numerical method [17] that is used to generate the reference solution here. Specifically, we compare the computational time and the memory usage using two representative cases, i.e., Kn $=0.1$ and 10.
We first discuss the computational time for MC-PINNs. As illustrated in Table 5, the computational time of MC-PINNs (second column in Table 5) is about 7 and 1.3 times that in the numerical method [17] for Kn $=0.1$ and 10, respectively. In the deep learning community, the transfer learning technique is a widely used approach to speed up the training of DNNs. Here we can also use it to speedup the training of MC-PINNs. Specifically, we first train the MC-PINNs from scratch for Kn $=1$, and then save the hyperparameters, i.e., weights and biases, in order to reuse them in other cases. In particular, we initialize the MC-PINNs for Kn $=0.1$ and 10 using the hyperparameters from the pre-trained MC-PINNs for Kn $=1$ and further train them for an additional number of steps until the convergence is reached. The relative $L_{2}$ errors for these two cases with transfer learning are $2.40\%$ and $2.6\%$, which are similar to the cases without transfer learning. Also, as we can see in Table 5, we can achieve about 3-6 times speedup with transfer learning, and the computational times for these two cases are now comparable to the numerical method in [17].
Note that we run the numerical method and MC-PINNs on the CPU and GPU, respectively. In particular, all the CPU computations are conducted on a single Intel Xeon Platinum 9242 processor, and the GPU computations are performed using an NVIDIA GeForce RTX 3090. We remark that the comparison present in Table 5 is not fair since we can also accelerate the numerical method by developing GPU codes and running the tests on GPUs. However, developing GPU codes for the numerical method is more challenging than MC-PINNs and is beyond the scope of the current study. We show the results here to provide a preliminary reference on the computational cost of MC-PINNs and one of the state-of-the-art numerical methods in this field. Therefore, we only present the computational time for the numerical method on CPUs.
@@
We now discuss the memory usage for MC-PINNs and the numerical method, which is present in Table 6. Note that the MC-PINNs with and without transfer learning have the same memory cost. We therefore only illustrate the memory usage in the MC-PINNs without transfer learning in Table 6. In the MC-PINNs, we employ the same number of training points for Kn $=0.1$ and 10, and thus the memory usage is almost the same for these two cases. In the numerical method, the memory cost is more expensive for Kn $=10$ than for Kn $=0.1$ since we have more discrete points in the angular domain for the former case. Also, the MC-PINN requires about $66\%$ and $16\%$ of the memory cost in the numerical method [17] for Kn $=0.1$ and 10, respectively. It is worth mentioning that the problem we consider here is relatively small. For real-world applications, we may have more than 40 meshes in each direction of the spatial domain in numerical methods. For example, if we increase the number of meshes in each direction from 40 to 400 and keep the same discretization in the solid angular space for Kn $=10$, the memory usage will increase to about 1,500GB, which is quite challenging to handle. However, in the MC-PINNs, we can use the same setup as used in the present case, suggesting that the memory usage can be the same, i.e., which is about 6,000 times less than in the numerical method [17].
@@
We would like to briefly summarize our main findings here: (1) the two-step sampling approach enables the MC-PINNs to use much less memory compared to the numerical method [17], especially for large-scale heat conduction problems spanning the diffusive and ballistic regimes; and (2) the computational time of MC-PINNs (GPU time) with transfer learning is comparable to the numerical method [17] (CPU time) for the specific problem considered here.